# Magnetic field controlled nucleation and size selection of silver nanoparticles

### v1

@@ Abstract
We examine the reduction of silver nanoparticle (AgNP) size under an external magnetic field within a classical nucleation theory framework combined with a sphere–packing description of atomic assembly. The model incorporates magnetic free-energy contributions arising from the coupling between the applied field and the magnetic susceptibility of the nucleating material, yielding a closed-form relation between nanoparticle radius and field strength. Our approach reproduces the experimentally observed decrease in the most-probable particle radius from approximately $170\,\mathrm{nm}$ at $\mathcal{B}=49.27\,\mathrm{mT}$ when the magnetic field is oriented parallel to the stirring plane, and to $\sim 155\,\mathrm{nm}$ at $\mathcal{B}=180.78\,\mathrm{mT}$ in the perpendicular configuration. Across the investigated field range, the theoretical predictions remain consistent with experimental measurements obtained under continuous mechanical stirring, supporting the interpretation that the observed size reduction originates from a magnetic-field-induced modification of the nucleation free-energy landscape. Within the limits of classical capillarity and spherical demagnetization, the results provide a physically transparent and computationally efficient framework for understanding magnetic-field-controlled nanoparticle size selection.
@@
Nanoparticles (NPs) exhibit chemical, optical and physical properties that differ markedly from their bulk counterparts, largely due to quantum confinement effects and their high surface-to-volume ratio [10.1016/j.jallcom.2023.172999, 10.1016/j.matpr.2023.05.456, doi.org/10.1016/j.ceramint.2024.06.067, 10.1007/s11051-024-06119-8, 10.1016/j.totert.2023.100038, 10.1039/D5YA00118H, D1MA00538C]. These size-dependent properties play a central role in applications ranging from photonics and sensing, protection of water infrastructures [apostolakis2024photoacoustic, Cousin2022] to biomedicine and energy conversion [D4MA90092H, Fabrizio12, henini2011handbook, Dong:20, 11142662, Uppu2021]. Note that NPs are not likely to replace silicon photonics entirely but are an emerging area of research that can significantly enhance and extend the capabilities of existing silicon technology, enabling new functionalities not possible with bulk silicon [Zafar2025-dl, Zafarpol4, Zafarreview, Zafarpol6, zafar2023band, zafar2023compact, ZafarIEEEAccess].

### v2

@@ Abstract
We examine the reduction of silver nanoparticle (AgNP) size under an external magnetic field within a classical nucleation theory framework combined with a sphere-packing description of atomic assembly. The model incorporates magnetic free-energy contributions arising from the coupling between the applied field and the magnetic susceptibility of the nucleating material, yielding a closed-form relation between nanoparticle radius and field strength. Our approach reproduces the experimentally observed decrease in the most-probable particle radius from approximately $170\text{ nm}$ at $\mathcal{B}=49.27\text{ mT}$ when the magnetic field is oriented parallel to the stirring plane, and to $155\text{ nm}$ at $\mathcal{B}=180.78\text{ mT}$ in the perpendicular configuration. Across the investigated field range, the theoretical predictions remain consistent with experimental measurements obtained under continuous mechanical stirring, supporting the interpretation that the observed size reduction originates from a magnetic-field-induced modification of the nucleation free-energy landscape. Within the limits of classical capillarity and spherical demagnetization, the results provide a physically transparent and computationally efficient framework for understanding magnetic-field-controlled nanoparticle size selection.
@@
Nanoparticles (NPs) exhibit chemical, optical and physical properties that differ markedly from their bulk counterparts, largely due to quantum confinement effects and their high surface-to-volume ratio [10.1016/j.jallcom.2023.172999, 10.1016/j.matpr.2023.05.456, doi.org/10.1016/j.ceramint.2024.06.067, 10.1007/s11051-024-06119-8, 10.1016/j.totert.2023.100038, 10.1039/D5YA00118H, D1MA00538C]. These size-dependent properties play a central role in applications ranging from photonics and sensing, protection of water infrastructures [apostolakis2024photoacoustic, Cousin2022] to biomedicine and energy conversion [D4MA90092H, Fabrizio12, henini2011handbook, Dong:20, 11142662, Uppu2021]. Note that NPs are not likely to replace silicon photonics entirely but are an emerging area of research that can significantly enhance and extend the capabilities of existing silicon technology, enabling new functionalities not possible with bulk silicon [Zafar2025-dl, Zafarpol4, Zafarreview, Zafarpol6, zafar2023band, zafar2023compact, ZafarIEEEAccess, Zafar:22]. Furthermore, NPs can also potentially replace more complex and expensive superlattice and multiple quantum structures for a wealth of optolectronic devices.[Pereira2020-ii, Vaks2022-uk, ma11010002, Michler1998-cp, Grempel].